the work function of silver is 4.73eV EM radiation with a frequency of 1.2x10^15 Hz strikes a piece of pure silver,what is the speed of the electrons that are emitted
1
Expert's answer
2020-04-28T09:34:59-0400

The photoelectric effect equation says

"hf=\phi+\frac{mv^2}{2}"

Hence, the speed of the electrons that are emitted

"v=\sqrt{\frac{2(hf-\phi)}{m}}"

"=\sqrt{\frac{2(6.67\times 10^{-34}\times 1.2\times 10^{15}-4.73\times 1.6\times 10^{-19})}{9.1\times 10^{-31}}}"

"=3.1\times 10^{5}\:\rm m/s"